OpenStudy (anonymous):

Does the following infinite geometric series diverge or converge? Explain. 7 + 21 + 63 + 189 + . .

OpenStudy (anonymous):

An infinite geometric series converges if its common ratio r satisfies –1 < r < 1. Otherwise it diverges.

OpenStudy (anonymous):

It converges; it has a sum. It converges; it does not have a sum. It diverges; it has a sum. It diverges; it does not have a sum.

OpenStudy (anonymous):

that means....r should be decimal

OpenStudy (anonymous):

it converges and has a sum diverges and has no sum

OpenStudy (anonymous):

I can only pick one :o

OpenStudy (mathmate):

Hint: \(common~ratio\) means a given term divided by the previous term. @PrincessKhiayla

OpenStudy (anonymous):

converges and have a sum

OpenStudy (anonymous):

Thank you @mathmate @lalithavasanth

OpenStudy (anonymous):

\[a _{n}\rightarrow \infty ~as~n \rightarrow \infty \] hence it does not converge.

OpenStudy (anonymous):

I'm confused :o

OpenStudy (anonymous):

\[a _{n}must \rightarrow 0~as~ n \rightarrow \]

OpenStudy (anonymous):

\[n \rightarrow \infty \]

OpenStudy (anonymous):

sorry i am leaving.

OpenStudy (anonymous):

So it does converge?

OpenStudy (mathmate):

@PrincessKhiayla First, you recognize that this is a geometric series since each term divided by the previous is a constant. This ratio is called the common ratio. Can you tell me the value of the common ratio r?

OpenStudy (anonymous):

It doesn't say @mathmate

OpenStudy (mathmate):

No it doesn't. You have to calculated it according to what I explained.

OpenStudy (anonymous):

Oh okay